# CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno

### v2

@@ Abstract.
Test generation tools attempt to automate the otherwise labor-intensive task of test creation,
but generating high-coverage
tests remains a challenge.
This paper proposes CoverUp, a novel approach to driving the
generation of high-coverage Python regression tests. CoverUp
iteratively improves test coverage, interleaving coverage analysis with
dialogs with the LLM that steer it to refine tests so that they increase coverage of
lines and branches.
We evaluate our prototype CoverUp implementation
across a benchmark of challenging
code derived from open-source Python projects, and show that
CoverUp substantially improves on the state of the art.
Compared to CodaMosa, a hybrid search/LLM-based test generator,
CoverUp achieves a per-module median line+branch
coverage of 80% (vs. 47%).
Compared to MuTAP, a mutation/LLM-based test generator,
CoverUp achieves an overall line+branch coverage of 90% (vs. 77%).
We show that CoverUp’s iterative, coverage-guided approach is
crucial to its effectiveness, contributing to nearly 40% of its successes.
@@
††ccs: Software and its engineering Software testing and debugging††footnotetext: Work done at the University of Massachusetts Amherst.
@@ 1. Introduction
Testing is essential to ensuring the quality of software, but
manually crafting tests can be so labor intensive that developers
choose not write them (Daka and Fraser, 2014).
Test generation tools attempt to automate this task, generally working
under the assumption that the source code is correct.
By generating tests based on it, they enable regression testing,
with the goal to prevent future bugs as the software is modified.
As these tools add tests that cover wider range of scenarios and execution paths,
they can help uncover bugs missed by manually written tests.
This paper proposes CoverUp, a novel approach to test generation
aimed aimed at achieving high coverage.
Our key insight is that large language models (LLMs) are capable of
simultaneously reasoning about code and coverage information.
@@
CoverUp incorporates detailed
coverage information into prompts customized to the current state
of the test suite, focusing the LLM on code that lacks coverage. Once
the LLM generates a set of tests, CoverUp executes them and
measures the resulting coverage. If the LLM-generated tests fail to
sufficiently improve coverage or fail to run altogether, CoverUp
continues the dialogue with the LLM, requesting changes to improve
coverage or fix the error.
In doing so, CoverUp in effect further refines and clarifies the
prompt, resulting in tests that significantly improve coverage.
Our empirical evaluation compares CoverUp to two state of the art
test generation systems: CodaMosa (Lemieux et al., 2023), a hybrid
search/LLM-based test generator, and MuTAP (Dakhel et al., 2024), a
a mutation/LLM-based approach.
We show that CoverUp substantially improves upon the state of the art.
Across a benchmark of challenging code derived from
open-source Python projects and using OpenAI’s GPT-4o LLM,
CoverUp increases coverage on every metric vs. CodaMosa, achieving
per-module median line+branch coverage of 80% (vs. 47%) and
overall line+branch coverage of 60% (vs. 45%).
Compared to MuTAP on one of the benchmark suites it supports,
CoverUp again achieves greater or equal coverage on every metric,
with an overall line+branch coverage of 90% (vs. 77%).
@@
It presents CoverUp, a novel approach that drives the
generation of high-coverage Python regression tests via a combination
of coverage analysis and large-language models;
It conducts a empirical analysis of our CoverUp prototype, showing that it
significantly advances the state of the art.
@@ 2. Related Work
Among the various proposed methods are
specification-based (Boyapati et al., 2002),
random (Miller et al., 1990; Fioraldi et al., 2023),
feedback-directed random (Pacheco et al., 2007),
symbolic execution guided random (“concolic”) (Godefroid et al., 2005; Sen et al., 2005; Tillmann and de Halleux, 2008),
search-based software testing (SBST) (Fraser and Zeller, 2012; Fraser and Arcuri, 2011; Lukasczyk and Fraser, 2022; Panichella et al., 2015, 2018)
and transformer-based approaches (Tufano et al., 2020).
The success of large language models on various tasks has motivated
their application to software engineering;
Wang et al. survey 102 recent papers using LLMs for software testing (Wang et al., 2024).
@@
Pynguin (Lukasczyk and Fraser, 2022) employs a search-based software testing (SBST)
approach. Starting from randomly created test cases, SBST employs genetic
algorithms to mutate the tests in an attempt to increase coverage.
Unfortunately, its search process can get stuck as test mutations repeatedly
lead to the same execution paths.
CodaMosa (Lemieux et al., 2023) addresses this problem by tracking
Pynguin’s progress; when it concludes that the search process has
stalled, it prompts an LLM for a test. It then uses that test to
re-seed the SBST, allowing it to resume progress.
@@
Bareiß et al. study the performance of the Codex LLM on
Java test case generation, among other code generation tasks (Bareiß et al., 2022).
Its prompts contain the signature of the method under test, an example of test generation,
and the body of the method; it discards any tests that do not compile.
By contrast, CoverUp’s prompts are based on code segments lacking coverage and
they explicitly request tests to improve it.
CoverUp also continues the chat with the LLM in case of build failure, failing tests,
or lack of coverage.
As Section 4.4 shows, this iterative dialogue is
responsibe for almost 50% of CoverUp’s successful test generation.
Vikram et al. discuss prompting LLMs based on API documentation
to generate property-based tests for Python (Claessen and Hughes, 2000; Vikram et al., 2023).
CoverUp instead bases its prompting on the source code and
coverage measurements and, while it accepts property-based tests
if the LLM generates them, it does not explicitly request them.
TiCoder prompts LLMs to generate tests based on
a natural language description of the intended functionality of code (Lahiri et al., 2022).
Rather than facilitate regression testing, however, TiCoder
generates tests with the goal of clarifying and formalizing user
intent.
TestPilot prompts an LLM to generate JavaScript unit tests based on the
function under test’s implementation, its documentation and usage snippets (Schäfer et al., 2024).
Like CoverUp, TestPilot checks the tests generated by the LLM and continues the chat
to refine the prompt in case of errors, but it does not prompt based on coverage, nor does it
continue the chat if the new tests do not improve coverage, nor does it use tool functions
to provide the LLM with additional context.
@@
Its use of LLMs is limited to prompting with code and for repairs when
a generated test fails compilation.
Unlike CoverUp, ChatUniTest does not employ coverage measurements
to indicate what lines or branches lack coverage, nor does it request
improvements if the generated tests do not improve coverage.
@@
The approaches below were developed concurrently with CoverUp, which was initially
posted on GitHub on August 7, 2023:
@@
The distillation LLM takes in arbitrary user input, such as documentation
and code examples, and generates prompts for the generation LLM, which
generates test inputs.
After the initial user input distillation and prompt selection, Fuzz4All repeatedly
employs the generation LLM, mutating its prompt to produce additional test inputs.
While Fuzz4All and CoverUp both use LLMs to generate test cases, their approaches
differ fundamentally: Fuzz4All generates test inputs based on documentation or examples,
relying on an user-provided test oracle for testing; CoverUp instead works based on
the source code and coverage measurements, looking to add tests that improve the test
suite’s coverage.
SymPrompt (Ryan et al., 2024) extracts approximate path constraints from Python functions
through static analysis and prompts an LLM for tests based on those constraints.
In contrast, CoverUp prompts an LLM directly indicating the lines and branches that
lack coverage.
Whereas SymPrompt extracts related type declarations and other code and include
these statically in its prompts, CoverUp provides a tool function (OpenAI, 2023)
that looks up arbitrary additional information about types, functions and other symbols,
thus allowing the LLM to drive this process.
SymPrompt’s error handling is limited to deleting lines in the response in an attempt
to correct syntax errors; CoverUp continues the chat in case of errors, asking the
LLM for a correction.
@@
SymPrompt is evaluated on an undisclosed subset of 897 functions (from at least 4,546)
previously used by CodaMosa and BugsInPy (Widyasari et al., 2020).
It achieves 74% line coverage using OpenAI’s GPT-4 LLM, approximately
a 2x improvement over its baseline prompt (Ryan et al., 2024).
Without knowing the exact functions employed despite multiple requests for
code and/or a replication package111Personal communication with SymPrompt authors.,
we can only provide a coarse comparison:
as Section 4.2 shows, using GPT-4o CoverUp
obtains 82% median per-module line coverage and 64% overall line coverage
on suite CM.
The code in that suite was also derived from CodaMosa’s evaluation, but contains
4,116 functions.
Even though CoverUp is evaluated on an 5x larger suite, the median per-module
coverage it achieves is also approximately a 2x improvement over an ablated version
of itself (Section 4.2).
TestGen-LLM (Alshahwan et al., 2024) is a test generation tool developed and deployed at
Meta to improve Kotlin unit tests.
TestGen-LLM prompts LLMs for tests that improve coverage, including in the prompt the
class under test and in some cases the existing test class.
It discards any tests that do not build, do not pass, or do not improve coverage, and
presents any remaining tests to a developer for approval during code review.
Like TestGen-LLM, CoverUp rejects tests that fail or lack coverage, but instead of
simply discarding them, it continues the chat with the LLM, asking for improvements.
CoverUp also indicates in its prompts what portions of the code lack coverage, and
allow the LLM to discover additional context through a tool function.
MuTAP (Dakhel et al., 2024) prompts an LLM
for Python unit tests based on mutation testing.
MuTAP first prompts for a test for a portion of code and, using it,
performs mutation testing.
It then prompts for new assertions for each surviving mutant, adding
these to the test.
Like CoverUp, MuTAP also prompts the LLM for a repair in case of syntax errors,
but, unlike CoverUp, not for other kinds of errors, lack of coverage,
nor does it include coverage information in its prompts.
Unfortunately, MuTAP currently lacks any ability to provide context for the code
under test and hardcodes the datasets used for evaluation in the code,
hampering efforts to evaluate it on other code.
In contrast, our implementation of CoverUp, like Pynguin and CodaMosa,
can be applied to any Python package.
@@
After measuring the test suite’s coverage, CoverUp prompts the
LLM so as to focus it on code segments that lack coverage.
It checks each new test, either accepting it if it increases coverage
or continuing the dialogue if it needs improvements.
CoverUp provides a tool function with the prompts with which
the LLM can request additional context.
@@
Given its incremental nature, once done CoverUp can be run again
to refocus the LLM on any segments missed in previous passes.
@@ 3. Approach
CoverUp first measures the code coverage obtained by the existing test suite and
uses it to identify code segments (functions and/or classes) that require additional testing
(Section 3.1).
It then prompts an LLM for tests for each of these segments,
combining information from coverage analysis and static code analysis
(Section 3.2).
CoverUp execute the LLM generated tests, once again measuring coverage.
If the tests do not compile, fail to run or increase coverage, CoverUp continues the
chat with prompts that request improvements and include any error messages.
(Section 3.3).
In both initial and continued prompts, CoverUp offers the LLM a tool function with
which it can request additional contextual information about names in the source code,
such as function or type definitions
(Section 3.4).
Having handled all code segments, CoverUp checks the extended test suite,
looking for any integration problems
(Section 3.5).
The entire process can be repeated an arbitrary number of times to allow the LLM
to re-focus on segments not covered by previous iterations.
@@ 3.1. Code Segmentation
It then uses the AST of each source file missing coverage to identify code segments
that need additional coverage.
Code segments typically consist of a single function or method; subject to a size limit,
however, they may contain an entire class.
As Figure 2 shows, if a segment contains a method,
CoverUp also includes a few lines of the original
class definition to provide the LLM with context, but omits other methods
and definitions.
Since we find that retaining comments improves code comprehension, CoverUp
does not remove these, thus trading a loss in concision for an increase in generation quality.
The goal is to provide to the LLM, with each prompt, a code excerpt
that is intelligible, provides enough context, includes the lines or branches
lacking coverage, and is as short as possible.
Keeping code segments short is important primarily because of LLMs’ limits on
their context (input) window.
While various current LLMs such as OpenAI’s GPT-4 and GPT-4o, Meta’s LLaMA 3 and
Anthropic’s Claude 3 support large context windows, of over 100K tokens,
this limit applies not only to the initial prompt, but to the total number of tokens
in the entire sequence of prompts and responses in the case of tool function calls
or a continued chat.
Even when prompts fit in these long context windows, shorter prompts remain preferable:
Rosas et al. (Rosas et al., 2024) show that
as the size of prompts containing code increases, so does the likelihood of inaccuracies.
Finally, LLM API use is often charged at a per-token rate, so that it literally
pays to be succinct.
Concretely, to identify the code segments in a source file, CoverUp first
computes a set of “interesting” lines: these are lines that either lack
coverage, or that are the source or the destination in branches that lack coverage.
For each of these lines, it then looks in the AST for a class, function or method
object containing that line.
If the object found is a class and it spans more lines than a configurable limit,
CoverUp adds that class definition as segment context and recursively looks
for another, smaller object containing the line.
@@
CoverUp also generates import statements that provide context
for each code segment: we find that without these, LLMs often make incorrect
assumptions about symbols in the code, leading to errors.
Rather than extract these verbatim from source code, CoverUp generates imports
based on static analysis.
In particular, it converts any relative imports to absolute imports: using relative
imports in code excerpts leads to these being used in tests outside of package
scope, where they are not valid.
We implement this step using SlipCover (Altmayer Pizzorno and Berger, 2023), a recently introduced
coverage analyzer that has near-zero overhead, and Python’s ast module.
@@
Figure 2. CoverUp summarizes method excerpts:CoverUp generates extremely compact code excerpts.
In this case, 200 lines of source code were originally
present between listing lines 1 and 2.

### v3

@@
Test generation tools attempt to automate the otherwise labor-intensive task of test creation, but generating high-coverage tests remains challenging.
This paper proposes CoverUp, a novel approach to driving the generation of high-coverage Python regression tests.
CoverUp combines coverage analysis, code context, and feedback in prompts that iteratively guide the LLM to generate tests that improve line and branch coverage.
We evaluate our prototype CoverUp implementation across a benchmark of challenging code derived from open-source Python projects and show that CoverUp substantially improves on the state of the art.
Compared to CodaMosa, a hybrid search/LLM-based test generator, CoverUp achieves a per-module median line+branch coverage of 80% (vs. 47%).
Compared to MuTAP, a mutation- and LLM-based test generator, CoverUp achieves an overall line+branch coverage of 90% (vs. 77%).
We also demonstrate that CoverUp’s performance stems not only from the LLM used but from the combined effectiveness of its components.
@@
††ccs: Computing methodologies Artificial intelligence††ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Testing is essential to ensuring software quality, but manually crafting tests can be so labor-intensive that developers choose not to write them (Daka and Fraser, 2014).
Test generation tools attempt to automate this task, generally assuming that the source code is correct.
By generating tests based on it, they enable regression testing, which aims to prevent future bugs as the software is modified.
As these tools add tests that cover a wider range of scenarios and execution paths, they can help uncover bugs missed by manually written tests.
This paper proposes CoverUp, a novel approach to test generation aimed at achieving high coverage.
Our key insight is that large language models (LLMs) can simultaneously reason about code and coverage information.
@@
CoverUp incorporates detailed coverage information into prompts customized to the current state of the test suite, focusing the LLM on code that lacks coverage.
Additionally, it provides a tool function (OpenAI, 2023) that allows the LLM to request additional source code context.
Once the LLM generates a set of tests, CoverUp executes them and measures the resulting coverage.
If the LLM-generated tests fail to improve coverage or fail to run altogether, CoverUp continues the dialogue with the LLM, requesting changes to improve coverage or fix the error.
In doing so, CoverUp, in effect, further refines and clarifies the prompt, resulting in tests that significantly improve coverage.
Our empirical evaluation compares CoverUp to two state-of-the-art test generation systems: CodaMosa (Lemieux et al., 2023), a hybrid search/LLM-based test generator, and MuTAP (Dakhel et al., 2024), a mutation/LLM-based approach.
We show that CoverUp substantially improves the state of the art.
Across a benchmark of challenging code derived from open-source Python projects and using OpenAI’s GPT-4o LLM, CoverUp increases coverage on every metric vs. CodaMosa, achieving per-module median line+branch coverage of 80% (vs. 47%) and overall line+branch coverage of 60% (vs. 45%).
Compared to MuTAP on one of the benchmark suites it supports, CoverUp again achieves greater or equal coverage on every metric, with an overall line+branch coverage of 90% (vs. 77%).
@@
It presents CoverUp, a novel approach that drives the generation of high-coverage Python regression tests via a combination of coverage analysis and large language models;
It conducts an empirical analysis of our CoverUp prototype, showing that it significantly advances the state of the art;
It conducts an ablation study, showing that CoverUp’s approach plays a substantial role in the prototype’s performance beyond what can be attributed to the LLM.
@@ 2. Related Work
Among the various proposed methods are specification-based (Boyapati et al., 2002), random (Miller et al., 1990; Fioraldi et al., 2023), feedback-directed random (Pacheco et al., 2007), symbolic execution guided random (“concolic”) (Godefroid et al., 2005; Sen et al., 2005; Tillmann and de Halleux, 2008), search-based software testing (SBST) (Fraser and Zeller, 2012; Fraser and Arcuri, 2011; Lukasczyk and Fraser, 2022; Panichella et al., 2015, 2018) and transformer-based approaches (Tufano et al., 2020).
The success of large language models on various tasks has motivated their application to software engineering; Wang et al. survey 102 recent papers using LLMs for software testing (Wang et al., 2024).
@@
Pynguin (Lukasczyk and Fraser, 2022) employs a search-based software testing (SBST) approach.
Starting from randomly created test cases, SBST employs genetic algorithms to mutate the tests, aiming to increase coverage.
Unfortunately, its search process can get stuck as test mutations repeatedly lead to the same execution paths.
CodaMosa (Lemieux et al., 2023) addresses this problem by tracking Pynguin’s progress; when it concludes that the search process has stalled, it prompts an LLM for a test.
It then uses that test to re-seed the SBST, allowing it to resume progress.
@@
Bareiß et al. study the performance of the Codex LLM on Java test case generation, among other code generation tasks (Bareiß et al., 2022).
Its prompts contain the signature of the method under test, an example of test generation, and the method’s body; it discards any tests that do not compile.
By contrast, CoverUp’s prompts are based on code segments lacking coverage, and they explicitly request tests to improve it.
CoverUp also continues the chat with the LLM in case of build failure, failing tests, or lack of coverage.
Section 4.4 shows that this iterative dialogue is responsible for nearly 40% of CoverUp’s successful test generation.
Vikram et al. discuss prompting LLMs based on API documentation to generate property-based tests for Python (Vikram et al., 2023). CoverUp instead bases its prompting on the source code and coverage measurements, and while it accepts property-based tests if the LLM generates them, it does not explicitly request them.
TiCoder prompts LLMs to generate tests based on a natural language description of the intended functionality of code (Lahiri et al., 2022).
Rather than facilitate regression testing, however, TiCoder generates tests to clarify and formalize user intent.
TestPilot prompts an LLM to generate JavaScript unit tests based on the function under test’s implementation, its documentation, and usage snippets (Schäfer et al., 2024).
Like CoverUp, TestPilot checks the tests generated by the LLM and continues the chat to refine the prompt in case of errors, but it does not prompt based on coverage, nor does it continue the chat if the new tests do not improve coverage, nor does it use tool functions to provide the LLM with additional context.
@@
Its use of LLMs is limited to prompting with code and for repairs when a generated test fails compilation.
Unlike CoverUp, ChatUniTest does not employ coverage measurements to indicate what lines or branches lack coverage, nor does it request improvements if the generated tests do not improve coverage.
@@
The approaches below were developed concurrently with CoverUp, which was initially posted on GitHub on August 7, 2023:
@@
The distillation LLM takes in arbitrary user input, such as documentation and code examples, and generates prompts for the generation LLM, which produces test inputs.
After the initial user input distillation and prompt selection, Fuzz4All repeatedly employs the generation LLM, mutating its prompt to produce additional test inputs.
While Fuzz4All and CoverUp both use LLMs to generate test cases, their approaches differ fundamentally: Fuzz4All generates test inputs based on documentation or examples, relying on a user-provided test oracle for testing; CoverUp instead works based on the source code and coverage measurements, looking to add tests that improve the test suite’s coverage.
SymPrompt (Ryan et al., 2024) focuses on generating tests that cover hard-to-reach code.
To do so, it first determines the path constraints needed to reach a certain part of the code and then prompts the LLM for tests based on those constraints.
CoverUp’s tests can also cover such code, but by prompting based on coverage measurements, a simpler and more direct approach.
While SymPrompt extracts related declarations and includes them statically in its prompts, CoverUp combines statically generated import declarations with a tool function that enables the LLM to drive the code context discovery process.
SymPrompt’s error handling is limited to deleting lines in the response in an attempt to correct syntax errors; CoverUp continues the chat in case of errors, asking the LLM for a correction.
@@
SymPrompt is evaluated on an undisclosed subset of 897 functions (from at least 4,546) previously used by CodaMosa and BugsInPy (Widyasari et al., 2020).
It achieves 74% line coverage using OpenAI’s GPT-4 LLM, approximately a 2x improvement over its baseline prompt (Ryan et al., 2024).
Without knowing the exact functions employed despite multiple requests for code and/or a replication package111Personal communication with SymPrompt authors., we can only provide a coarse comparison: as Section 4.2 shows, using GPT-4o CoverUp obtains 82% median per-module line coverage and 64% overall line coverage on a 4,116-function suite also derived from CodaMosa’s evaluation.
Even though CoverUp is evaluated on a 5x larger suite, the median per-module coverage it achieves is also approximately a 2x improvement over an ablated version of itself (Section 4.2).
TestGen-LLM (Alshahwan et al., 2024) is a test generation tool developed and deployed at Meta to improve Kotlin unit tests.
TestGen-LLM prompts LLMs for tests that improve coverage, including in the prompt the class under test and, in some cases, the existing test class.
It discards any tests that do not build, do not pass, or do not improve coverage and presents any remaining tests to a developer for approval during code review.
Like TestGen-LLM, CoverUp rejects tests that fail or lack coverage, but instead of simply discarding them, it continues the chat with the LLM, asking for improvements.
CoverUp also indicates in its prompts what portions of the code lack coverage and allows the LLM to discover additional context through a tool function.
MuTAP (Dakhel et al., 2024) prompts an LLM for Python unit tests based on mutation testing.
MuTAP first prompts for a test for a portion of code and, using it, performs mutation testing.
It then prompts for new assertions for each surviving mutant, adding these to the test.
Like CoverUp, MuTAP also prompts the LLM for a repair in case of syntax errors, but unlike CoverUp, it does not prompt it for improvements in case of other errors or lack of coverage, nor does it include coverage information in its prompts.
Unfortunately, MuTAP currently lacks any ability to provide context for the code under test and hardcodes the datasets used for evaluation in the code, hampering efforts to evaluate it on other code.
In contrast, our implementation of CoverUp, like Pynguin and CodaMosa, can be applied to any Python package.
@@
After measuring the test suite’s coverage, CoverUp prompts the LLM, focusing it on code segments that lack coverage.
It checks each new test, either accepting it if it increases coverage or continuing the dialogue if it needs improvements.
CoverUp provides a tool function with which the LLM can request additional context.
@@
Given its incremental nature, once done, CoverUp can be rerun to refocus the LLM on any segments missed in previous passes.
@@ 3. Approach
CoverUp first measures the code coverage obtained by the existing test suite and uses it to identify code segments (functions and/or classes) that require additional testing (Section 3.1).
It then prompts an LLM for tests for each segment, combining information from coverage analysis and static code analysis (Section 3.2).
CoverUp next executes the LLM-generated tests, once again measuring coverage.
If the tests do not compile, fail to run, or do not increase coverage, CoverUp continues the chat with prompts that request improvements and include any error messages (Section 3.3).
In both initial and continued prompts, CoverUp offers the LLM a tool function with which it can request additional contextual information about names in the source code, such as function or type definitions (Section 3.4).
Having handled all code segments, CoverUp checks the extended test suite, looking for any integration problems (Section 3.5).
The entire process can be repeated arbitrarily, enabling the LLM to refocus on segments not covered in previous iterations.
@@ 3.1. Code Segmentation
It then uses the AST of each source file missing coverage to identify code segments that need additional coverage.
Code segments typically consist of a single function or method; subject to a size limit, however, they may contain an entire class.
As Figure 2 shows, if a segment contains a method, CoverUp also includes a few lines of the original class definition to provide the LLM with context but omits other methods and definitions.
We find that retaining comments improves code comprehension; for that reason, CoverUp does not remove them, thus trading a loss in concision for an increase in generation quality.
The goal is to provide the LLM, with each prompt, a code excerpt that is intelligible, provides enough context, includes the lines or branches lacking coverage, and is as short as possible.
Keeping code segments short is important primarily because of LLMs’ limits on their context (input) window.
While various current LLMs, such as OpenAI’s GPT-4 and GPT-4o, Meta’s LLaMA 3, and Anthropic’s Claude 3, support large context windows of over 100K tokens, this limit applies to more than just the initial prompt.
It applies to the number of tokens in the entire sequence of messages included with each request, which grows with each tool function call, LLM response, and continued chat prompt.
Even when prompts fit in these long context windows, shorter prompts remain preferable: Rosas et al. (Rosas et al., 2024) show that as the size of prompts containing code increases, so does the likelihood of inaccuracies.
Finally, LLM providers typically charge on a per-token basis, so it literally pays to be succinct.
Concretely, to identify the code segments in a source file, CoverUp first computes a set of “interesting” lines: these are lines that either lack coverage or are the source or the destination in branches that lack coverage.
It then looks in the AST for a class, function, or method object containing each line.
If the object found is a class that spans more lines than a configurable limit, CoverUp adds that class definition as segment context and recursively looks for another, smaller object containing the line.
@@
We implement this step using SlipCover (Altmayer Pizzorno and Berger, 2023), a recently introduced coverage analyzer with near-zero overhead, and Python’s ast module.
@@
Figure 2. CoverUp summarizes method excerpts:CoverUp generates compact code excerpts.
In this case, 200 lines of source code were originally present between listing lines 1 and 2.